OpenStudy (anonymous):

how do you solve (y-3)(-y+8)=9y

OpenStudy (anonymous):

expand the brackets: -y^2 + 8y +3y -24 - 9y = 0 -y^2 + 2y -24 = 0 y^2 - 2y + 24 = 0 y = 2 + sqrt(4 - 4*1*24) / 2 = 2 +-sqrt (-92) / 2 = 1 + sqrt23 i or 1 - sqrt23 i